BCG Status Updates

May 29, 2014 - WIMR Electrical Shutdown (again)

On May 29th starting at 8pm Boldt will be cutting the power to everything, including the red emergency power outlets. All fume hoods, BSC's and outlets will be off during this shutdown. The power will be shut off from 8pm until 10pm and then turned back on from 10pm until midnight. See previous entry for details.

May 28, 2014 - WIMR Electrical Shutdown

On May 28th starting at 8pm Boldt will be cutting the power to everything, including the red emergency power outlets. All fume hoods, BSC's and outlets will be off during this shutdown. The power will be shut off from 8pm until 10pm and then turned back on from 10pm until midnight, to allow freezers and incubators to recover operational temps, and then turned off again for 2 more hours until all power is restored at 2am.

Boldt was made aware of concerns regarding loss of temp in our freezers and incubators and we expect that by keeping the power off for no more than two hour increments we will not see critical loss of temp in any of our equipment. However, cold rooms will not be restored to operational temp until the following morning at 6am.

-Do not plan to be in the building after 8 pm on the 28th as the building will go dark and all access points will fail locked

-Do not leave processes running over night as there will be no power

-Please find another option for items stored in cold rooms that can't tolerate room temp for 8 hours.

-Please post contact information on all of your freezers to help us reach out to you if we discover a problem with one of your units.

-Shutdown any PC attached to an instrument to avoid a hard shutdown.

SMPH Facilities Mangagment

Tags: 

April 23, 2014 - Wireless Issue

10:20am: the SMPH wireless network is currently acting erratically.

10:35am: issue resolved, but a permanent fix may require additional downtime.

March 21, 2014 - Outage Day

  • 12:15pm - outage complete; all compute servers are back on.
  • 11:15am - chilled water is flowing again; we've started to turn back on the compute servers.
  • 10:10am - temperature still stable; chilled water refill expected to start at 11:00am.
  • 7:30am - server room temperature stabilized.
  • 6:45am - chilled water turned off.
  • 6:30am - steamfitters on site.

Scheduled Friday AM Outage Update

We're still on schedule for a Friday AM outage. The steamfitters have the parts they need to do the maintenance. They have assigned a steamfitter to just our server room during the outage, to make sure we have everything we need to respond to cooling issues.

Scheduled Outage: Friday, March 21, 2014

To make repairs on the chilled water system, the Physical Plant will be shutting down chilled water for several hours Friday morning, starting at about 7:00am.

Chilled water is essential for the correct operation of Biostatistics' main server room in the CSC. When the water stops flowing, the temperature shoots up to 100-110F in about 20-30 minutes.

We believe we will be able to maintain core services during the chilled water outage. This includes file services (over NFS or Samba) and the departmental web pages. However, research computational machines will be shut off. This includes private hosts, as well as a large number of servers that are used by condor.

The physical plant has been working with us closely to make sure we are able to preserve this basic functionality during this maintenance. However, there's always the possibility of something going wrong. We will keep people informed as we learn more. The weather on Friday could cause this maintenance to be rescheduled.

We will shut down compute-only hosts Thursday evening, to be ready for the Physical Plant the next morning.
We will likely shut down Condor Friday AM.
We will notify owners of private compute hosts which of their machines will be shut off.
We will update this web page with minor details and updates, and will email everyone if there are major service changes.

Please send email to sysreq@biostat.wisc.edu with questions.

Pages